OpenStudy (anonymous):

How do you solve 2^x=3^x? It is a exponential equation.

OpenStudy (anonymous):

you can see right away that x=0 works

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

What about e^t=50?

OpenStudy (anonymous):

If the exponent is a variable, it is an exponential equation.

OpenStudy (anonymous):

ln 2^x= ln 3^x x ln 2= x ln 3 x(ln 2- ln3)=0

OpenStudy (amistre64):

ln(2) - ln (3) = ln(2/3) ;)

OpenStudy (amistre64):

at any rate: x(C) = 0 means x = 0

OpenStudy (anonymous):

satellite not here to shout EYEBALL

OpenStudy (anonymous):

imran, can u help with other one too?

OpenStudy (anonymous):

sure